OpenStudy (anonymous):

According to the Kinetic Molecular Theory, as the volume of an enclosed gas increases: the number of particle impacts per unit area increases the temperature must change the total number of particle impacts increase the number of particle impacts per unit area decreases

OpenStudy (photon336):

\[pV = nRT\] let's keep pressure constant, and number of moles constant. nR = constant \[V_{1} = T_{1} nR \] \[\frac{ V_{1} }{ T_{1} } = nR = \frac{ V_{2} }{ T_{2} }\] \[\frac{ V_{1} }{ T_{1} } = \frac{ V_{2} }{ T_{2} }\] If we plot volume and temperature we get a straight line.  if the volume goes up the temperature goes up too.

OpenStudy (photon336):

To test the last claim we need to keep the temperature constant to study pressure and volume. \[pV = nRT \] \[P_{1}V_{1} = nR = k = P_{2}V_{2}\] IF increase the volume or make V_2 bigger, for the two to equal echoer, the pressure must go down. pressure translates to how many times the particles hit the walls of the container fewer times. so that's also true.

OpenStudy (anonymous):

thank you sir much obliged

OpenStudy (anonymous):

Select the TWO variables that are held constant when testing Boyle's Law in a manometer. pressure temperature volume of gas mass of gas atmospheric pressure?

OpenStudy (anonymous):

i cant decide which one to pick

OpenStudy (photon336):

@23188 no problem, I provided you an explanation for your other question too.